OpenStudy (anonymous):

Find how long it will take a $5000 investment to grow to $10,000 if it is invested at 8% interest compounded quarterly.. a=p(1+r/n)^nt

OpenStudy (anonymous):

Take the time as x. Make both the bases equal and solve it.

OpenStudy (anonymous):

\[10,000=5,000(1+\frac{.08}{4})^{4t}\] \[2=(1.002)^{4t}\] \[4t=\frac{\ln(2)}{\ln(1.002)}\] \[t=\frac{\ln(2)}{4\ln(1.002)}\] then a calculator

OpenStudy (anonymous):

how did you do step 3? i know ln, but how did the2 move over?

OpenStudy (anonymous):

step 1: divide by 5,000 step 2: you solve \[b^x=A\] for x via \[x=\frac{\ln(A)}{\ln(b)}\] by the "change of base formula"

OpenStudy (anonymous):

step 3: divide by 4

OpenStudy (anonymous):

Thanks.